# SafeVLA: Towards Safety Alignment of Vision-Language-Action Model via Safe Reinforcement Learning

### v1

@@ Abstract
Vision-language-action models (VLAs) have shown great potential as generalist robot policies. However, these models pose urgent safety challenges during deployment, including the risk of physical harm to the environment, the robot itself, and humans. How can safety be explicitly incorporated into VLAs? In this work, we propose SafeVLA, a novel algorithm designed to integrate safety into VLAs, ensuring the protection of the environment, robot hardware and humans in real-world settings. SafeVLA effectively balances safety and task performance by employing large-scale constrained learning within simulated environments. We demonstrate that SafeVLA outperforms the current state-of-the-art method in both safety and task performance, achieving average improvements of 83.58% and 3.85%, respectively, in simulation. By prioritizing safety, our approach eliminates high-risk behaviors and reduces the upper bound of unsafe behaviors to 1/35 of that in the current state-of-the-art, thereby significantly mitigating long-tail risks. Furthermore, the learned safety constraints generalize to diverse, unseen scenarios, including multiple out-of-distribution perturbations and tasks. Our data, models and newly proposed benchmark environment are available at https://sites.google.com/view/pku-safevla.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: The overview of SafeVLA pipeline. Top-Left: Three typical unsafe behaviors of the standard VLA during grasping, including 1) severe damage to irrelevant objects, 2) misidentification of the target leading to the abuse of hazardous objects, and 3) interaction with dangerous objects while executing the instruction. Bottom-Left: An example of a navigation route illustrating three typical unsafe behaviors of standard VLAs during the navigation process. Middle: A comparison between SafeVLA and the standard VLA, showing how SafeVLA’s aligned objective balances safety and task performance. Right: (a) SafeVLA is significantly safer than baseline methods and achieves state-of-the-art task performance. (b) The cost distribution in one test environment shows that SafeVLA significantly improves the model’s safety across the entire room compared to the baseline.
@@
Leveraging large-scale pre-training datasets (Kaplan et al., 2020) and advanced post-training techniques (Ouyang et al., 2022), foundation models have achieved remarkable success (Bommasani et al., 2021; Firoozi et al., 2023; Zhou et al., 2024a). Specifically, vision-language models (VLMs) (OpenAI, 2023) expand on large language models (LLMs) by understanding both text and images, allowing them to follow multimodal instructions (Alayrac et al., 2022; Liu et al., 2024b). Building on this, vision-language-action models (VLAs) (Brohan et al., 2022; O’Neill et al., 2023; Team et al., 2024; Kim et al., 2024) take it a step further, allowing robots to follow vision-language instructions and perform tasks in real-world environments. As VLAs continue to evolve, they have the potential to develop into generalist robot policies (Reed et al., 2022; Ma et al., 2024), that can execute previously unseen instructions and effectively generalize behaviors across a wide range of robot embodiments, scenes, skills, and objects (O’Neill et al., 2023). In view of the increasing complexity and power of these foundation models (Dubey et al., 2024; OpenAI, 2024a; Liu et al., 2024a), ensuring their alignment with human values and safety has become more critical than ever before (Kaddour et al., 2023; Ji et al., 2023).
One line of these works has documented various safety risks in LLMs and VLMs. These risks include generating misleading content (Chen & Shu, 2023), producing discriminatory statements (Gallegos et al., 2024), facilitating criminal activities (Caldwell et al., 2020; Dalrymple et al., 2024), and amplifying biases (Henderson et al., 2023), which has sparked widespread concern (Hendrycks et al., 2023; Shevlane et al., 2023; Tu et al., 2023; Bailey et al., 2023; Ji et al., 2024b; Zhao et al., 2024). Another line of these works, focuses on ensuring the safety alignment of LLMs and VLMs by proposing strategies such as data augmentation (Gulcehre et al., 2023), content moderation (Inan et al., 2023; Chi et al., 2024), reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Touvron et al., 2023), and lightweight alignment (Ji et al., 2024a).
However, the urgent safety challenges posed by VLAs remain unresolved. Since these models can interact directly with the physical world, ensuring the safety of their generated robot behaviors becomes significantly more complex (Guiochet et al., 2017). Beyond concerns tied to the text modality alone, robots executing tasks must address physical risks, including potential harm to the environment, the robot’s hardware, and humans. In addition, mental effects, such as sensory and psychological disturbances in nearby humans, must also be considered (Vicentini, 2020). For instance, when a robot performs tasks in a room, it must not damage the environment or itself, nor change the state of unrelated objects. When interacting with humans, the robot should limit its speed and torque to avoid rushing the task, as this could lead to offense or distress. Despite large-scale multitask behavior cloning and careful alignment in existing VLAs (Hu et al., 2024; Zhang et al., 2024), even the most advanced models have yet to explicitly define and integrate safety as an integral aspect of their design (Brohan et al., 2023; Gu et al., 2023; Ehsani et al., 2024; Belkhale et al., 2024; Black et al., 2024; Liu et al., 2024c). This fundamental limitation introduces safety flaws within the model itself, limiting the deployment of VLAs in real-world applications (Zacharaki et al., 2020; Falco et al., 2021).
To address this challenge, we propose SafeVLA, an algorithm that incorporates safety into VLAs while optimizing task performance. Specifically, safety is learned through large-scale constrained learning, using contrasting safety-positive and safety-negative samples within simulated environments. SafeVLA explicitly formalizes the safety constraints of VLAs in real-world scenarios via the constrained Markov decision process (CMDP) paradigm (Altman, 2021; Ji et al., 2024d), ensuring that the pre-trained VLA model adheres to safety requirements and, within these constraints, maximizes task performance. Our approach achieves state-of-the-art results in both safety and task performance, yielding average improvements of 83.58% and 3.85% over the current best method (Hu et al., 2024), respectively, demonstrating its effectiveness in balancing these objectives. Furthermore, quantitative evaluation shows that SafeVLA trained models generalize well to unseen scenarios that span multiple out-of-distribution (OOD) perturbations (e.g., lighting, material, and color variations).
To the best of our knowledge, SafeVLA is the first algorithm that explicitly incorporates safety constraints into VLAs. We summarize our contributions as follows:
  • We introduce Safety-CHORES, a new simulation benchmark with millions of unique scenes that incorporate safety constraints. The diversity of these scenes ensures that the safety constraints introduced during training are scalable and generalizable, while also enabling a comprehensive evaluation of model safety. Our qualitative and quantitative analyses expose inherent shortcomings in current VLAs: their failure to integrate safety considerations explicitly.
  • We propose the SafeVLA framework, which directly addresses the critical limitation of existing VLAs by incorporating explicit safety modeling through a constrained learning paradigm.
  • We demonstrate through a series of experiments: from simple to complex tasks, that SafeVLA achieves state-of-the-art results in both safety and task performance. The performance and safety gains of our approach generalize across 12 distinct OOD task combinations, derived from 4 perturbations and 3 tasks, across over 500 procedurally generated scenes.
  • We will open-source our code, data, models, and the newly proposed benchmark environment to facilitate further research in this area.
@@ 2 Related Work
Vision-Language-Action Models
Recently, end-to-end vision-language-action models (VLAs), which directly predict robot actions, have made rapid progress. This is due to breakthroughs in computer vision (Awais et al., 2025) and natural language processing (Zhou et al., 2024a). Notably, these advances not only improve individual model components but also enable a more seamless integration of vision, language, and action. VLAs typically use foundation models from these fields as core components (Brohan et al., 2023) and are trained on large-scale robot trajectory datasets that encompass a wide range of robot embodiments, scenes, skills, and objects (O’Neill et al., 2023). These models are capable of following multi-task natural language instructions for robot tasks such as manipulation (e.g., RT-1 (Brohan et al., 2022), RT-2 (Brohan et al., 2023), RT-X (O’Neill et al., 2023), OpenVLA (Kim et al., 2024), Octo (Team et al., 2024), $\pi_{0}$ (Black et al., 2024), RDT(Liu et al., 2024c)) and navigation (e.g., ViNT (Shah et al., 2023)), generalizing well to unseen settings and outperforming expressive imitation learning methods built from scratch, such as Diffusion Policy (Chi et al., 2023). Existing research has focused on expanding the training scale and model size (Brohan et al., 2022, 2023), utilizing existing VLMs as the backbone (Zawalski et al., 2024), adjusting architectures to enhance model capabilities (Black et al., 2024; Liu et al., 2024c; Pertsch et al., 2025), and improving the robustness and generalizability of these models (Wang et al., 2024).
In this work, we select SPOC (Ehsani et al., 2024) as our initial model for alignment. SPOC is a transformer-based model pre-trained on millions of frames of expert trajectories. It is specifically designed to follow natural language instructions in an unbounded number of procedurally generated household scenes (Deitke et al., 2022). These tasks require coordinating both navigation and manipulation in complex scenarios. This challenge exceeds the capabilities of models that focus solely on manipulation or navigation, allowing a more comprehensive evaluation of the safety of robot behavior.
Safety Alignment
AI alignment aims to ensure that the behavior of AI systems is in line with human intentions and values (Ji et al., 2023). Foundation models pre-trained on massive datasets have already demonstrated their remarkable capability and versatility across various domains. Examples include the CLIP series (Radford et al., 2021) for image-language understanding, the SAM series (Kirillov et al., 2023) for image segmentation, the GPT series (OpenAI, 2023) for text and vision-text dialogue, and Sora model (OpenAI, 2024b) for video generation. However, the outputs of these models sometimes lead to unforeseen negative consequences, such as assisting in criminal activities (Caldwell et al., 2020; Dalrymple et al., 2024), generating misinformation (Chen & Shu, 2023), amplifying biases (Henderson et al., 2023), and producing harmful content (Ganguli et al., 2022). Therefore, ensuring the alignment of the behavior of these models is critically important (Amodei et al., 2016).
Extensive prior work has explored methods for aligning foundation models, particularly language and vision-language models, with safety objectives. RLHF-based methods (Ouyang et al., 2022; Bai et al., 2022) align models with binary human preferences for generated content, yet they often struggle to balance complex trade-offs between safety and task performance. To address this, some research has explored the modeling of different evaluation dimensions using separate models (Touvron et al., 2023). This strategy helps mitigate the challenge, though it still relies heavily on hyperparameter tuning. Safe-RLHF further refines it by optimizing a Lagrangian dual objective (Dai et al., 2023), ensuring that safety constraints are prioritized. Furthermore, some researchers (Zhou et al., 2024b; Ji et al., 2024c) propose alignment objectives based on language feedback, offering richer supervision for model-generated content.
In the context of aligning VLAs, existing work focuses primarily on fine-tuning these models using reinforcement learning (RL) to improve generalization ability and performance, rather than aligning their behavior with human intentions and values. The settings in FLaRe (Hu et al., 2024) and GRAPE (Zhang et al., 2024) are most similar to ours, as both begin with a pre-trained VLA model and use RL to align the model with task objectives. However, FLaRe does not address safety concerns and GRAPE is limited to fixed desktop manipulation scenarios. Their primary goal is to enable the policy to generalize to more diverse settings, thereby enhancing task performance. In contrast, SafeVLA explicitly incorporates safety within a constrained learning framework, ensuring that safety constraints are prioritized and satisfied.
@@ 3 Problem Formulation
We formulate robotic control tasks as the language-conditioned constrained Markov decision process (CMDP) (Altman, 2021) $(\mathcal{S},\mathcal{A},r,\mathcal{C},\mathbb{P},\mathcal{L},\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $r$ the reward function, $\mathcal{C}$ the constraint set, $\mathbb{P}$ the transition probability function, $\mathcal{L}$ the set of natural language instructions, $\mu_{0}$ the initial state distribution, and $\gamma$ the discount factor. Following the setup in FLaRe (Hu et al., 2024), the reward function $r$ is defined as a sparse reward function that outputs a binary value based on a natural language instruction $l\in\mathcal{L}$ and the current state $s_{t}\in\mathcal{S}$, indicating whether the task has been successful.
The constraint set $\mathcal{C}=\left\{(c_{i},b_{i})\right\}_{i=1}^{m}$ consists of cost functions $c_{i}$ and corresponding cost thresholds $b_{i}$. Each cost function $c_{i}$ produces a positive discrete value based on the single-step state transition from $s_{t}$ to $s_{t+1}$ and the action $a_{t}$. This value quantifies the magnitude of the safety cost incurred for the constraint $i$ at the time step $t$. We define the reward return as
@@
$\mathcal{J}(\pi_{\theta})=\mathbb{E}_{\pi_{\theta},\mathcal{L}}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(l,s_{t}\right)\right],$ (1)
which represents the expected cumulative reward when following a policy $\pi_{\theta}$ for a given instruction $l\in\mathcal{L}$. Similarly, the cost return for constraint $i$ is given by

### v2

# SafeVLA: Towards Safety Alignment of Vision- Language-Action Model via Constrained Learning

@@ Abstract
Vision-language-action models (VLAs) show potential as generalist robot policies. However, these models pose extreme safety challenges during real-world deployment, including the risk of harm to the environment, the robot itself, and humans. How can safety constraints be explicitly integrated into VLAs? We address this by exploring an integrated safety approach (ISA), systematically modeling safety requirements, then actively eliciting diverse unsafe behaviors, effectively constraining VLA policies via safe reinforcement learning, and rigorously assuring their safety through targeted evaluations. Leveraging the constrained Markov decision process (CMDP) paradigm, ISA optimizes VLAs from a min-max perspective against elicited safety risks. Thus, policies aligned through this comprehensive approach achieve the following key features: (I) effective safety-performance trade-offs, this exploration yields an 83.58% safety improvement compared to the current state-of-the-art method, while also maintaining task performance (+3.85%). (II) strong safety assurance, with the ability to mitigate long-tail risks and handle extreme failure scenarios. (III) robust generalization of learned safety behaviors to various out-of-distribution perturbations. Our data, models and newly proposed benchmark environment are available at https://pku-safevla.github.io.
@@ 1 Introduction
Figure 1: The Integrated Safety Approach (ISA) pipeline. Our proposed pipeline employs multi-faceted framework for the systematic safety alignment of vision-language-action (VLA) models.
@@
Embodied AI aims to develop a generalist policy that can perform perception, interaction, reasoning, and adaptation in the physical world liu2024aligning . Building on the emergence of large language models (LLMs) and vision-language models (VLMs), vision-language-action models (VLAs) brohan2022rt ; o2023open ; team2024octo ; kim2024openvla advance this field by enabling robots to follow vision-language instructions and perform tasks in real-world environments. As these models continue to evolve, they have the potential to become generalist robot policies reed2022generalist ; ma2024survey , capable of executing previously unseen instructions and effectively generalizing behaviors across a diverse range of robot embodiments, scenes, skills, and objects o2023open . Ensuring the alignment of these models with human values and safety has become more critical than ever kaddour2023challenges ; ji2023ai , due to their increasing complexity and power dubey2024llama ; openai2024o1 ; liu2024deepseek . While significant progress has been made in task performance, the explicit integration of safety mechanisms remains an open challenge.
How can safety constraints be explicitly integrated into VLAs without loss of performance?
The safety risks of LLMs and VLMs have been extensively studied, with existing methods such as data augmentation gulcehre2023reinforced , content moderation inan2023llama ; chi2024llama , reinforcement learning from human feedback (RLHF) ouyang2022training ; touvron2023llama , Safe-RLHF dai2023safe , and lightweight alignment ji2024aligner . However, these safety mechanisms cannot be directly applied to VLAs, as there is a substantial gap between the abstract safety concerns at the model intention level and the unique safety challenges posed by the complex and unpredictable physical world guiochet2017safety . Despite large-scale behavior cloning and careful alignment in existing VLAs hu2024flare ; zhang2024grape , the most advanced models have yet to explicitly define and integrate safety as an integral aspect of their design brohan2023rt ; gu2023rt ; ehsani2024spoc ; belkhale2024rt ; black2024pi0 ; liu2024rdt . This fundamental limitation motivates an urgent need to explore methodologies capable of explicitly embedding safety constraints into the VLAs zacharaki2020safety ; falco2021governing .
To tackle this challenge, we make the first systematic explorations into VLA safety alignment. Our approach is grounded in the constrained Markov decision process (CMDP) framework altman2021constrained ; ji2024omnisafe , leveraging methodology from safe reinforcement learning (SafeRL) for optimization. We investigate an integrated safety approach (ISA), which systematically considers four key aspects: comprehensively modeling safety requirements within the CMDP setup, actively eliciting diverse unsafe behaviors to inform constraints, rigorously constraining VLA policies using CMDP-compliant SafeRL techniques, and thoroughly assuring safety through targeted evaluations. The core insight of such an approach is to explicitly trade off safety and task performance, prioritizing safety adherence. Our investigation addresses the significant engineering challenges in adapting and scaling these principles for VLAs, focusing on how to effectively model, elicit, and utilize safety signals.
To the best of our knowledge, this work is the first systematic explorations into explicitly integrating safety constraints into VLAs using principles from SafeRL. Our main contributions are:
  • Integrated Safety Approach (ISA) Exploration: We conduct a comprehensive investigation into an ISA for VLA safety alignment. This involves systematically exploring and implementing methodologies for: (a) modeling intricate safety requirements and diverse scenarios; (b) eliciting a wide spectrum of latent unsafe behaviors; (c) constraining VLA policies using CMDP-based SafeRL, optimizing from a min-max perspective; and (d) establishing robust practices for assuring the safety of aligned policies through targeted evaluations and stress-testing. Our study details how these interconnected aspects contribute to a more holistic safety alignment.
  • Environment: Addressing the gap in comprehensive VLA safety assessment, we introduce Safety-CHORES. This novel testbed is a direct result of the modeling and eliciting aspects of our ISA. To this end, the benchmark is designed with fine-grained safety constraints embedded within diverse, long-horizon tasks that integrate navigation and manipulation. By incorporating large-scale procedurally generated scenes and specifically targeting safety critical components, Safety-CHORES more effectively surfaces VLA vulnerabilities than conventional benchmarks.
  • Empirical Validation and Key Findings: Our extensive experiments demonstrate that policies aligned through our ISA exploration achieve: (I) an effective trade-off between safety and task performance, evidenced by an average 83.58% safety improvement over state-of-the-art method, while maintaining task performance (+3.85%); (II) strong safety assurance, particularly in mitigating long-tail risks and handling extreme failure scenarios, as supported by the elimination of high-risk actions and a drastic reduction in unsafe incident severity; and (III) robust generalization of learned safety behaviors to out-of-distribution (OOD) perturbations. These findings underscore the potential of a comprehensive, multi-faceted approach to significantly advance VLA safety.
@@ 2 Related Work
Vision-Language-Action Models.  Vision-language-action models (VLAs) brohan2022rt ; brohan2023rt ; o2023open ; ehsani2024spoc ; team2024octo ; kim2024openvla ; black2024pi0 represent a significant step towards generalist robots capable of executing complex tasks based on multimodal instructions in diverse environments reed2022generalist ; ma2024survey . These models, often built upon powerful foundation models zawalski2024robotic ; liu2024rdt ; pertsch2025fast and trained on large-scale trajectory datasets o2023open , demonstrate impressive task performance and generalization ability wang2024towards . However, their real-world deployment is hindered by safety concerns inherent to physical interaction guiochet2017safety ; zacharaki2020safety . While safety alignment is actively researched for LLMs and VLMs ouyang2022training ; bai2022constitutional ; dai2023safe ; ji2024beavertails ; ji2024aligner , methods focusing on mitigating abstract risks like harmful content generation ganguli2022red ; hendrycks2023overview do not readily address the concrete physical hazards faced by embodied agents. Current VLA training, typically relying on imitation learning (IL) ehsani2024spoc or standard reinforcement learning (RL) fine-tuning hu2024flare ; zhang2024grape , lacks mechanisms for explicitly integrating and enforcing safety constraints, leaving a critical gap for reliable deployment falco2021governing .
Safe Reinforcement Learning.  Safe reinforcement learning (SafeRL) within the constrained Markov decision process (CMDP) framework altman2021constrained ; gu2022review , offers a formal approach to policy optimization where an agent learns to maximize task rewards while explicitly satisfying predefined safety constraints. This paradigm contrasts with heuristic methods like reward shaping, which indirectly encode safety preferences and lack formal guarantees. While SafeRL techniques have been explored for aligning foundation models (e.g., Safe-RLHF dai2023safe ), applying them to high-dimensional, multimodal VLAs operating in complex physical environments poses unique challenges act2024artificial . Model-free, first-order optimization methods compatible with the CMDP formulation, such as Lagrangian-based approaches PIDLag2020 ; dai2023augmented , are promising for VLAs as they avoid restrictive assumptions about system dynamics or state structure, making them suitable for learning from raw perceptual inputs like RGB images ji2024omnisafe . Our work systematically explores the application of these principles to VLA safety alignment.
Benchmarking Safety and VLA Alignment.  Evaluating VLA safety requires appropriate benchmarks capable of eliciting unsafe behaviors. Existing SafeRL benchmarks often involve simplified dynamics or non-photorealistic settings leike2017ai ; yuan2022safe ; ji2023safety ; tomilin2025hasard , while standard VLA benchmarks primarily focus on task success across manipulation james2020rlbench ; mees2022calvin ; zhang2024vlabench or navigation anderson2018vision ; deitke2020robothor , lacking diverse and challenging scenarios with built-in safety constraints. Thus, we propose Safety-CHORES to comprehensively assess safety alongside task performance in complex, procedurally generated environments. While prior work like FLaRe hu2024flare and GRAPE zhang2024grape employed RL fine-tuning for VLAs, their objective was primarily task performance improvement and generalization, without the explicit safety constraint satisfaction central to our SafeRL-based approach. Our approach utilizes the CMDP framework to formulate VLA alignment as a constrained optimization problem. This approach differs fundamentally from prior RL fine-tuning methods. Specifically, it allows for directly tackling the trade-off between safety and task performance to ensure adherence to predefined safety constraints.
@@ 3 Problem Formulation
Constrained Markov Decision Process.  The constrained Markov decision process (CMDP) (altman2021constrained, ) is commonly used to model dynamic decision-making under uncertainty when multiple objectives are present. In this framework, the policy aims to maximize one objective while satisfying constraints on the others. A CMDP is defined as a tuple $(\mathcal{S},\mathcal{A},\mathbb{P},r,\mathcal{C},\mu,\gamma)$, where $\mathcal{S}$ is state space, $\mathcal{A}$ is action space. $\mathbb{P}(s^{{}^{\prime}}|s,a)$ is probability of state transition from $s$ to $s^{{}^{\prime}}$ after playing $a$. $r(\cdot):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and $r(s^{{}^{\prime}}|s,a)$ denotes the reward that the agent observes when state transition from $s$ to $s^{{}^{\prime}}$ after it plays $a$. The set $\mathcal{C}=\{(c_{i},b_{i})\}_{i=1}^{m}$, where $c_{i}$ are cost functions: $c_{i}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and limits are $b_{i}$, $i=1,\cdot,m$. $\mu(\cdot):\mathcal{S}\rightarrow[0,1]$ is the initial state distribution and $\gamma\in(0,1)$. Let $\mathcal{H}_{t}$ be the set of all possible trajectories $(s_{0},a_{0},\dots,s_{t-2},a_{t-2},s_{t-1})$ of length $t$.
From CMDP to VLA Safety Alignment.  To address the safety constrained decision-making problem in VLAs, we formulate VLA safety alignment using an adapted CMDP framework, defined by the tuple $(\mathcal{S},\mathcal{A},\mathbb{P},r,\mathcal{C},\mathcal{L},\mu,\gamma)$, where $\mathcal{L}$ is the set of natural language instructions. The reward function $r$ is conditioned on a natural language instruction $l\in\mathcal{L}$, and is defined as $r:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\times\mathcal{L}\rightarrow\mathbb{R}$. Let $\pi_{\bm{\theta}}$ denote the vision-language-action model parameterized by $\bm{\theta}$, which maps an observation history $h_{t}=(o_{t+1-H},a_{t+1-H},\dots,o_{t})$ to an action $a_{t}\sim\pi_{\bm{\theta}}(\cdot|l,h_{t})$, where $H>1$ is the temporal horizon, $l$ is the natural language instruction. Each observation $o_{t}=(v_{t},p_{t})$ represents the multimodal perceptual input at time $t$, comprising visual input $v_{t}$ and proprioceptive input $p_{t}$.
@@
The reward-return is defined as $\mathcal{J}(\pi_{\bm{\theta}})=\mathbb{E}_{\pi_{\bm{\theta}},\mathcal{L}}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(s_{t+1}|s_{t},a_{t},l\right)\right]$. The set of feasible policies is then defined as